# Semantic Refinement with LLMs for Graph Representations

### v1

@@ Appendix I Prompts
For each node, we build an initial description by concatenating (i) an original description and (ii) a topological summary. For citation graphs, the original description is the node text (e.g., paper title and abstract). For datasets without node text (e.g., airport graphs), this component is omitted. The topological summary encodes structural cues using a fixed natural-language schema: we first state the global graph context (graph type, node type, number of nodes, edge type, and number of edges), and then append node-level property statements where each property is reported with its scalar value and its rank among all nodes. The exact verbalization template is shown in Figure 5 and follows the consistent schema adapted from wang2024tans.
Iterative refinement.
Starting from the initialized description above, we refine the target node text using a single unified prompt wrapper. Only the placeholders (highlighted tokens such as GRAPH_TYPE) are swapped per node/dataset. Two blocks are optional: (i) a target history block that summarizes previous rewrite attempts and the GNN feedback signals (top probability and predictive entropy), and (ii) an example block that provides a small set of training nodes as reference for how descriptions behave under the GNN. Figure 6 shows the complete prompt layout.
@@
[Optional: Target history (most recent first)]
Iteration ITER_ID: GNN prediction = PRED_NAME (CORRECTNESS); top prob = TOP_PROB; entropy = ENTROPY_VAL.
@@
Figure 6: Unified prompt used to refine node text descriptions. Only placeholders (e.g., GRAPH_TYPE) are filled at runtime; the history and example blocks are optional.

### v2

@@ Appendix I Implementation Details
For each node, we build an initial description by concatenating (i) an original description and (ii) a topological summary. For citation graphs, the original description is the node text (e.g., paper title and abstract). For datasets without node text (e.g., airport graphs), this component is omitted. The topological summary encodes structural cues using a fixed natural-language schema: we first state the global graph context (graph type, node type, number of nodes, edge type, and number of edges), and then append node-level property statements where each property is reported with its scalar value and its rank among all nodes. The exact verbalization template is shown in Figure 6 and follows the consistent schema adapted from Wang et al. (2025a).
Semantic refinement.
Starting from the initialized description above, we refine the target node text using a single unified prompt wrapper. Only the placeholders (highlighted tokens such as GRAPH_TYPE) are swapped per node/dataset. An optional example block provides a small set of training nodes as reference for how descriptions behave under the GNN, serving as in-context calibration data. Figure 7 shows the complete prompt layout.
@@
[Optional: Target Predictive Performance]
GNN prediction = PRED_NAME (CORRECTNESS); top prob = TOP_PROB; entropy = ENTROPY_VAL.
@@
Figure 7: Unified prompt used to refine node text descriptions. Only placeholders (e.g., GRAPH_TYPE) are filled at runtime; the prediction and example blocks are optional.